0:00 Both the LBTI and LUCI WFS brought up. The SOUL OCAM had triangular noise pattern first try, but a cycle of the BCU resolved.

AZ Science

2:10 Slewing to first target HOP65, AOref magnitude R 9.8, Running ESM
  • OCAM mag 10.9, seeeing 1-1.2"
  • bin1 freq 519, slope rms 0.574
  • tt 10.8 ho1 11.39, optimize gain went a bit bananas on the gains, plots level at 10 and 10 so not crazy in the end
  • TN 20220312_021800, 022852, 023339 (reduced ho1 a touch), 032941
  • 0.55" on luci1 and 0.65" on luci2
3:33 Moving on to telluric, no ESM.

3:52 Preset to HOPS 163, Catalog magnitude R=12.79
  • OCAM not live during close loop (and no it was not just the normal reconfig). Interrupt command since the star looked like it was walking off. Reset binnings recovered. Resumed
  • Bin2, freq 400, seeing 1.5 on the DIMM but we are down at 1.6 airmass
  • LUCI did not recognize the preset done. It still says its waiting for collimation. Had to wait for luci to time out.
  • Alex resent the preset by accident so need to close loop again. I just manually ramped gains to approximately where the flat part of the curve was in the previous optimize gain. TT10, HO1 7
  • OCAM mag 14.5, bin2, freq 400
  • Elevation is 35, airmass 1.74 and setting
  • TN 042341, 042813
4:48 Done with LUCI AZ programs and handing over to ILocater. Safing the LUCI SOUL WFSbut leaving powered up in case eng time is handed over.

Ilocater

5:17 sending preset to lamda leo, catalog magnitude V4.31. Srtar not found. these positions baysides are very different from the last time we were on sky so star not easily found. Slewing to alpha leo.

5:45 There is some sort of vignetting on the ilocater pupil image. adjusting tertiary tip tilt. Absolute tilt tip:750, tilt 100
  • Alpha Leo bin1 900 Hz, Gain 4
  • there is an angle coming through the system so adjusting baysides.
  • New bayside positions 62.1, 35.8, 57.9
  • Image looks out of focus
6:20 Issue: The loop claims closed on WFS and adsec with supposed gains applied but loop is not closed in reality with target wandering off and no real gains on there. Just resending AO preset. We were opening and closing in rapid fired adjusting baysides

6:32 New z stage focus for baysides determined to be 64.4. now time to play with ncpa
  • Found Z7 -5, Z9 15 resulted in a pretty good psf, determied with optical gain around 0.3. saved to Ilocater_wam (yes wam...). Sym linked to LBTI in

    /home/sxwunit/live/aoroot/calib/wfs/current/ncpa

  • Seeing 0.95", optical loop data set with NCPA on 20220312_071737
7:40. Jonathan and team have noted a regular open loop fram every 10 seconds. the adsec mirror gui is showing the "pending skip frame" counter incrementing at about that cadence. Could this be NCPA related? It is not abovious on the OCAM or mirror gui but is visible on their scicam monitoring,
  • TN 074117
  • NCPA updates at a cadence of every 10seconds
  • Once we turned off NCPA there were no dips 075737
  • 1164|2022-03-12 07:45:57.783716| MAIN > arbFunc updateNcpaInternal (, 'LBTIWFS', 'LBTI', 0, 0)
    EG
    WfsArbitrator_29526|DEB| 1169|2022-03-12 07:46:17.837399| MAIN > arbFunc updateNcpaInternal done.
    WfsArbitrator_29526|DEB| 1170|2022-03-12 07:46:27.843431| MAIN > arbFunc updateNcpaInternal (, 'LBTIWFS', 'LBTI', 0, 0)
    WfsArbitrator_29526|DEB| 1171|2022-03-12 07:46:27.857353| MAIN > arbFunc updateNcpaInternal done.
    WfsArbitrator_29526|DEB| 1172|2022-03-12 07:46:37.861768| MAIN > arbFunc updateNcpaInternal (, 'LBTIWFS', 'LBTI', 0, 0)
    WfsArbitrator_29526|DEB| 1173|2022-03-12 07:46:37.876965| MAIN > arbFunc updateNcpaInternal done.
    WfsArbitrator_29526|DEB| 1174|2022-03-12 07:46:47.881450| MAIN > arbFunc updateNcpaInternal (, 'LBTIWFS', 'LBTI', 0, 0)
    WfsArbitrator_29526|DEB| 1175|2022-03-12 07:46:47.899845| MAIN > arbFunc updateNcpaInternal done.
    WfsArbitrator_29526|DEB| 1176|2022-03-12 07:46:57.904908| MAIN > arbFunc updateNcpaInternal (, 'LBTIWFS', 'LBTI', 0, 0)
    WfsArbitrator_29526|DEB| 1177|2022-03-12 07:46:57.920089| MAIN > arbFunc updateNcpaInternal done.
    WfsArbitrator_29526|DEB| 1178|2022-03-12 07:47:07.924380| MAIN > arbFunc updateNcpaInternal (, 'LBTIWFS', 'LBTI', 0, 0)
    WfsArbitrator_29526|DEB| 1179|2022-03-12 07:47:07.936690| MAIN > arbFunc updateNcpaInternal done.
    WfsArbitrator_29526|DEB| 1180|2022-03-12 07:47:17.941099| MAIN > arbFunc updateNcpaInternal (, 'LBTIWFS', 'LBTI', 0, 0)
    WfsArbitrator_29526|DEB| 1181|2022-03-12 07:47:17.954941| MAIN > arbFunc updateNcpaInternal done.
    WfsArbitrator_29526|DEB| 1182|2022-03-12 07:47:27.959690| MAIN > arbFunc updateNcpaInternal (, 'LBTIWFS', 'LBTI', 0, 0)
    WfsArbitrator_29526|DEB| 1183|2022-03-12 07:47:27.974395| MAIN > arbFunc updateNcpaInternal done.
    WfsArbitrator_29526|DEB| 1184|2022-03-12 07:47:37.981487| MAIN > arbFunc updateNcpaInternal (, 'LBTIWFS', 'LBTI', 0, 0)
    WfsArbitrator_29526|DEB| 1185|2022-03-12 07:47:37.996368| MAIN > arbFunc updateNcpaInternal done.
    WfsArbitrator_29526|DEB| 1186|2022-03-12 07:47:48.000515| MAIN > arbFunc updateNcpaInternal (, 'LBTIWFS', 'LBTI', 0, 0)
    WfsArbitrator_29526|DEB| 1187|2022-03-12 07:47:48.012682| MAIN > arbFunc updateNcpaInternal done.
    WfsArbitrator_29526|DEB| 1188|2022-03-12 07:47:58.018212| MAIN > arbFunc updateNcpaInternal (, 'LBTIWFS', 'LBTI', 0, 0)
8:04 sending preset to HD100518, V mag 6.55
  • OCAM magnitude 7.1. Bin1, freq 900, OCAM gain 262.
  • 1" seeing
  • No NCPA this time
  • companion nearby 0.1" away
9:22 HD117478 preset, Vmag 8.42
  • close binary
  • OCAM mag 8.8
  • bin1 freq 900 OCAM gain 518
  • closed with NCPA this time, seeing 0.82"
  • TN 083311

Back to AZ Science

Ilocater finished early and rather than scrabmle for LBTO eng tasks we are going to pick up AZ science in the good conditions.

8:48 LUCI2 MOS error (writing here just because)

Rotator position 83, elevation 58

Straing gauges: SG1 on, SG2 off, but it was a precondition violation stran gauge switch 1 and 2 active.

X clamp hold, closed MRC, Y clamp is hold closed MRC so mask 7 is in FPU. Picker arms open

Rotator 125.8 ok, spindle ok,

9:03 Preset to SpARCS AORef catalog mag 13.53
  • OCAM magnitude 15.2
  • Running bin2, freq 400, OCAM gain 600, 3 photons/subapp
  • gains 2.36, 3.66, 0.08 after optimization
11:55 Done AZ ESM science. Set safe for WFS.

-- JenniferPower - 12 Mar 2022
Topic revision: r3 - 13 Mar 2022, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback